# Defining implication relation for classical logic

### v11

@@ 4.1 Semantics of IRL
An interpretation of a formal language assigns a meaning (semantic value) to each expression within that language. However, logic is fundamentally concerned with the truth values of the statements in its language, rather than their specific content. So interpretations that assign the same truth value (even if they assign different contents) to a statement can be viewed as the same in a logic. Hence, an interpretation of a propositional language can be represented by just a truth assignment or valuation function.
Let $\{\mathrm{T,F}\}$ be the set of truth values for two-valued logics.
@@
Let $U=\{\mathrm{T,F}\}^{L}$, the set of all possible valuation functions for $L$.
The set $L$ is countably infinite (as the set of primitive propositions is finite or countably infinite) and $|U|=|\{\mathrm{T,F}\}|^{|L|}=2^{|L|}=2^{\aleph_{0}}=\mathfrak{c}$, so $U$ is generally uncountable. (In classical logic, interpretations of formulas are reduced to that of primitive propositions as all the operations are truth-functional, so there are only $2^{n}$ interpretations where $n$ is the number of primitive propositions.)
@@
The T-set of a formula $\phi$ is defined as
$V_{\phi}=\{v\in U\mid v(\phi)=T\}.$
In terms of T-set, the usual definitions of validity/tautology, semantic consequence and equivalence, are defined below.
  • Formulas $\phi$ and $\psi$ are semantically equivalent if $\phi\vDash\psi$ and $\psi\vDash\phi$, which is equivalent to $V_{\phi}=V_{\psi}$ by the definition of semantic consequence and the algebra of sets.
@@
The semantics of IRL is defined according to the meaning of operations, via T-set, as follows.
@@
Semantics of bi-implication is given by a collection of characterizing conditions as it is not truth-functional. Implication is defined by bi-implication.
More properties of the operations can be derived from the definition of the semantics. An example is of the tautological implication:
$V_{\phi\to\psi}=U\text{ if and only if }V_{\phi}\subseteq V_{\psi},$
since $V_{\phi\wedge\psi\leftrightarrow\phi}=U$ is equivalent to $V_{\phi\wedge\psi}=V_{\phi}$, $V_{\phi\wedge\psi}=V_{\phi}\cap V_{\psi}$, and $V_{\phi}\cap V_{\psi}=V_{\phi}$ is equivalent to $V_{\phi}\subseteq V_{\psi}$ with the algebra of sets. In terms of the original meaning of implication, this result indicates that $V_{\phi}\subseteq V_{\psi}$ is just the “certain mechanism” of the tautological implication from $\phi$ to $\psi$ (the scope of the mechanism is $U$).
Two special cases of $V_{\phi}\subseteq V_{\psi}$ are ϕ=italic-$\phi=\bot$ and $\psi=\top$. Since $\emptyset\subseteq V_{\psi}$ and $V_{\phi}\subseteq U$, we have $V_{\bot\to\psi}=U$ and $V_{\phi\to\top}=U$. This indicates that ECQ and its dual do not conflict with the original meaning of implication.
The algebraic system IRLA is developed (cf. Subsection 4.2) with a language more concise than that of the semantic system in T-set notation. Essentially, IRLA is just a representation of the semantic system obtained by replacing T-set notation with a more concise notation, as defined in Table 1.
Table 1: T-set and IRLA representation
@@
    The IRLA representations of the T-sets $x=V_{\phi}$, $y=V_{\psi}$ and $x\leftrightarrow y=V_{\phi\leftrightarrow\psi}$ imply the definition $V_{\phi}\leftrightarrow V_{\psi}=V_{\phi\leftrightarrow\psi}$, and similarly $V_{\phi}\to V_{\psi}=V_{\phi\to\psi}$.
@@ 4.2 Algebra IRLA
Let $X$ be the underlying set, $a,b,...$ constants, $x,y,...$ variables, and $s,t,...$ terms of the algebraic language.
@@
  • Axioms: commutativity, associativity and distributivity for $\wedge$ and $\vee$; identity elements for $1$ and 00; complement element for $\neg$. Where associativity is redundant (Huntington (1904)).
@@
Remark 4.2.1.
Although these rules are actually default for (a variety of) algebras such as Boolean algebra so that they are usually not listed, it is best to specify the set of logical rules especially for an algebraic system that represents a logical system (e.g. Boolean algebra and Boolean lattice), even if these rules may be used tacitly.
@@
A Boolean lattice $(X,1,0,\neg,\wedge,\vee,\leq)$ is an extention of Boolean algebra $(X,1,0,\neg,\wedge,\vee)$ by adding the definition of partial order ($\leq$): $x\leq y\Leftrightarrow x\wedge y=x$. The following can be derived:
@@
The algebraic system $IRLA(X,1,0,\neg,\wedge,\vee,\leftrightarrow,\to,\leq)$ is the extension of Boolean lattice $BL(X,1,0,\neg,\wedge,\vee,\leq)$ by defining the new operations ($\leftrightarrow$,$\to$ , →) with the axioms below:
@@ 5     $=x[x\mapsto y]\leftrightarrow x$ Selected replacement
6     $=y\leftrightarrow x$ Result of replacement
7 $x\leftrightarrow y\leq y\leftrightarrow x$ 1–6: Transitivity of $=$ and $\leq$
@@ 2     $\leq(x\leftrightarrow y)[y\mapsto z]$ Replacement property of $\leftrightarrow$ (Axiom 1)
3     $=(x\leftrightarrow z)$ Result of replacement
@@
Theorem 4.3.4 (Leibniz’s replacement for bi-implication in IRLA).
@@ 9 $x=y$ 5,8: Antisymmetry of $\leq$
Combined with $x=y\Rightarrow x\leftrightarrow y=1$ (Axiom 2) we get $x\leftrightarrow y=1\Leftrightarrow x=y$
@@
By definition of $\to$ (Axiom 3), $x\to y=1\Leftrightarrow x\wedge y\leftrightarrow x=1$, by definition of $\leq$, $x\leq y\Leftrightarrow x\wedge y=x$. In terms of the above theorem, $x\wedge y\leftrightarrow x=1\Leftrightarrow x\wedge y=x$, so $x\to y=1\Leftrightarrow x\leq y$. ∎
In terms of Theorem 4.3.5 and Corollary 4.3.1, an IRLA theorem of the form $s\leftrightarrow t=1$ or $s\to t=1$ follows immediately from a corresponding theorem of Boolean lattice (BL) of the form $s=t$ or $s\leq t$. Examples of such theorems of IRLA and BL are listed in Table 2.
Table 2: Theorems of IRLA and BL
Name Theorem of IRLA Theorem of BL
@@
Theorem 4.3.6 (Bi-implication and implication).
$x\leftrightarrow y=(x\to y)\wedge(y\to x).$
Proof.
$x\leftrightarrow y$
1     $=(x\leftrightarrow y)\wedge(x\to x)$ Reflexivity of $\to$
2     $\leq x\to y$ Replacement property of $\leftrightarrow$ (Axiom 1)
3 $x\leftrightarrow y\leq x\to y$ 1–2: Transitivity of $=$ and $\leq$
4 $x\leftrightarrow y\leq y\to x$ 1–3: similarly
5 $x\leftrightarrow y\leq(x\to y)\wedge(y\to x)$ 3,4: One-to-two of $\leq$
$(x\to y)\wedge(y\to x)$
6     $=(x\wedge y\leftrightarrow x)\wedge(x\wedge y\leftrightarrow y)$ Definition of $\to$
7     $\leq x\leftrightarrow y$ Transitivity of $\leftrightarrow$
8 $(x\to y)\wedge(y\to x)\leq x\leftrightarrow y$ 6–7: Transitivity of $=$ and $\leq$
9 $x\leftrightarrow y=(x\to y)\wedge(y\to x)$ 5,8 Antisymmetry of $\leq$
Corollary 4.3.2 (Bi-implication elimination in IRLA).
$x\leftrightarrow y\leq x\to y.$
Proof.
Because $x\leftrightarrow y=(x\to y)\wedge(y\to x)\leq x\to y$ by conjunction elimination. ∎
Theorem 4.3.7 (Modus ponens in IRLA).
$(x\to y)\wedge x\leq y.$
Proof.
$(x\to y)\wedge x$
1     $=(x\wedge y\leftrightarrow x)\wedge x$ Definition of $\to$
2     $\leq x\wedge y$ Replacement property of $\leftrightarrow$
3     $\leq y$ Conjunction elimination
4 $(x\to y)\wedge x\leq y$ Transitivity of $=$ and $\leq$
Theorem 4.3.8 (Antisymmetry of implication in IRLA).
@@ 4 $(x\to y)\wedge x\leq y$ Transitivity of $=$ and $\leq$
1     $=(x\wedge y\leftrightarrow x)\wedge(x\wedge y\leftrightarrow y)$ Definition of $\to$
@@ 1     $=(x\wedge y\leftrightarrow x)\wedge(y\wedge z\leftrightarrow y)$ Definition of $\to$
3     $\leq(x\wedge y\leftrightarrow x)\wedge(x\wedge(y\wedge z)\leftrightarrow x)$ Replacement of $\leftrightarrow$
@@ 4     $=(x\wedge y\leftrightarrow x)\wedge((x\wedge y)\wedge z\leftrightarrow x)$ Associativity
5     $\leq x\wedge z\leftrightarrow x$ Replacement of $\leftrightarrow$
@@ 1     $=(x\wedge z\leftrightarrow z)\wedge(y\wedge z\to z)$ Definition of $\to$
2     $\leq x\wedge y\wedge z\leftrightarrow z$ Replacement
@@ 4 $(z\to x)\wedge(z\to y)\leq z\to x\wedge y$ 1–3: Transitivity of $=$ and $\leq$
5     $=(z\to x\wedge y)\wedge(x\wedge y\to x)$ Conjunction elimination
@@ 10 $(z\to x)\wedge(z\to y)=z\to x\wedge y$ 4,9: Antisymmetry of $\leq$
Theorem 4.3.11 (Maximum of implication).
@@
1 $(x\to y)\wedge x\leq y$ Replacement property of $\leftrightarrow$
2 $x\to y\leq\neg x\vee y$ 1: Law in BL ($x\wedge y\leq z\Leftrightarrow x\leq\neg y\vee z$)
@@
It follows from $1\to 0\leq\neg 1\vee 0=0\vee 0=0$ by Theorem 4.3.11 and boundedness $0\leq 1\to 0$. ∎
@@ 4.4 Soundness and completeness
By definition (cf. Subsection 4.1), the T-set of IRL formula schema $\eta(\phi,\psi,...)$ is the IRLA term $t(x,y,...)$; and $\vDash\phi$ is the same as $x=1$, $\vDash\phi\leftrightarrow\psi$ is the same as $x=y$ or $x\leftrightarrow y=1$, and $\vDash\phi\to\psi$ is the same as $x\leq y$ or $x\to y=1$.
@@
For each axiom schema $\eta(\phi,\psi,...)$ of IRL, verify $t(x,y,...)=1$ in IRLA. For example, $\phi\leftrightarrow\phi$ is a tautology since $x\leftrightarrow x=1$ in IRLA. ∎
@@
The three primitive rules of IRL – Conjunction introduction $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$, Conjunction elimination $\phi\wedge\psi\Rightarrow\phi$ and Replacement property $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$, are tautologousness-preserving, that is, if their premises are tautologies, so are their conclusions.

### v12

@@ 6 $\phi_{1}\wedge\phi_{2}\to\psi$ 4,5: Transitivity
Meta Rule 3.4.1 (Equivalence of implication theorem and disjunction theorem).
$(\vdash\phi\to\psi)\Leftrightarrow(\vdash\neg\phi\vee\psi).$
Proof.
1 $\vdash\phi\to\psi$ Premise
2     $(\phi\to\psi)\to\neg\phi\vee\psi$ Implication-to-Disjunction
3     $\phi\to\psi$ 1 (assumed theorem)
4     $\neg\phi\vee\psi$ 2,3: Modus ponens
5 $\vdash\neg\phi\vee\psi$ 2–4: Unconditional proof
1 $\vdash\neg\phi\vee\psi$ Premise
2     $\phi$ Premise
3     $\neg\phi\vee\psi$ 1 (assumed theorem)
4     $\psi$ 2,3: Disjunctive syllogism
5 $\vdash\phi\to\psi$ 3–5: Conditional proof
@@ 4.1 Semantics of IRL
An interpretation of a formal language assigns a meaning (semantic value) to each expression within that language. However, logic is fundamentally concerned with truth values rather than specific contents of statements of a language. So interpretations that assign the same truth value (even if they assign different contents) to a statement can be viewed as the same in a logic. Hence, an interpretation of a propositional language can be represented by just a truth assignment or valuation function.
Let $\{\mathrm{T,F}\}$ be the set of truth values for a two-valued logic.
@@
Let $U=\{\mathrm{T,F}\}^{L}$, the set of all valuation functions for $L$.
The set $L$ is countably infinite (even in a case when the countable set of primitive propositions is finite), thus $|U|=|\{\mathrm{T,F}\}|^{|L|}=2^{|L|}=2^{\aleph_{0}}=\mathfrak{c}$, i.e. $U$ is generally uncountable with the cardinality of $\mathbb{R}$. (In classical logic, as all the operations – including particularly the implication – are truth-functional, the number of all interpretations for a formula with $n$ propositional variables is only $2^{n}$.)
@@
The T-set of a formula $\phi\in L$ is defined as
$V_{\phi}=\{v\in U\mid v(\phi)=\mathrm{T}\}.$
@@
The semantic system for IRL is defined via T-set as follows.
@@
    The structure $(2^{U},U,\emptyset,{}^{\complement},\cap,\cup,\subseteq)$ is the power set Boolean lattice of $U$. (For any formula $\phi$ its T-set $V_{\phi}$ is a member of $2^{U}$ by definition.)
  2. 2.
@@
Remark 4.1.1.
  • Semantics of bi-implication, as it is not truth-functional, is given by a collection of characterizing conditions. Implication is defined by bi-implication.
  • A derived result for implication is $V_{\phi\to\psi}=U\text{ if and only if }V_{\phi}\subseteq V_{\psi}$. In terms of the original meaning of implication, this indicates that $V_{\phi}\subseteq V_{\psi}$ is just the “certain mechanism” of the tautological implication from $\phi$ to $\psi$ (the scope of this mechanism is $U$).
  • By definition of T-set, for any formula $\phi$, $\emptyset\subseteq V_{\phi}\subseteq U$, thus $V_{\bot\to\psi}=U$ and $V_{\phi\to\top}=U$, indicating that ECQ and its dual have “certain mechanisms” of tautological implications, so they do not conflict with the original meaning of implication.
Definition 4.1.2 (Validity and semantic consequence).
With respect to the semantic system, validity/tautology, semantic consequence and equivalence, are defined below:
  • Formulas $\phi$ and $\psi$ are semantically equivalent if $\phi\vDash\psi$ and $\psi\vDash\phi$, or equivalently $V_{\phi}=V_{\psi}$.
The semantic system of IRL defined in this subsection can be viewed as an extension of the power set Boolean lattice of $U$, $(2^{U},U,\emptyset,{}^{\complement},\cap,\cup,\subseteq)$, by adding semantic characteristics of propositional formulas of IRL in the T-set notation. The algebraic system IRLA developed in the next subsection, an extension of the Boolean lattice $(X,1,0,\neg,\wedge,\vee,\leq)$, is simply a “concise version” of the semantic system via T-set in this subsection. Symbols of IRLA language are defined in Table 2.
Table 2: Definition of IRLA symbols
@@
    The IRLA representation of the T-sets $x=V_{\phi}$, $y=V_{\psi}$ and $x\leftrightarrow y=V_{\phi\leftrightarrow\psi}$ implies the definition $V_{\phi}\leftrightarrow V_{\psi}=V_{\phi\leftrightarrow\psi}$, and similarly $V_{\phi}\to V_{\psi}=V_{\phi\to\psi}$.
@@ 4.2 Algebra IRLA
Let $(X,1,0,\neg,\wedge,\vee)$ be an algebraic system, where $X$ is the underlying set, $(1,0,\neg,\wedge,\vee)$ are operations of arities $(0,0,1,2,2)$ on $X$. Note that “an operation on a set” implies that the set is closed under the operation. So $1,0\in X$, i.e. $X$ contains at least two constants $1$ and 00. And $X$ may contains other countably many constants $a,b,...$.
Let $x,y,...$ be variables and $s,t,...$ terms for the algebraic language.
@@
  • Axioms: commutativity, associativity and distributivity for $\wedge$ and $\vee$; definitions of identity elements $1$ and 00; definition of complement or negation $\neg$. Associativity is redundant (Huntington (1904)).
@@
Although the above rules are actually default for (a variety of) algebras such as Boolean algebra – so they are usually not listed explicitly, it is best to specify the set of logical rules especially for an algebraic system that represents a logical system (e.g. Boolean algebra and Boolean lattice), even if these rules may be used tacitly.
@@
A Boolean lattice $(X,1,0,\neg,\wedge,\vee,\leq)$ is an extension of Boolean algebra $(X,1,0,\neg,\wedge,\vee)$ by adding the defining axiom/rule for the partial order ($\leq$): $x\leq y\Leftrightarrow x\wedge y=x$. The following can be derived:
@@
The algebraic system $IRLA=(X,1,0,\neg,\wedge,\vee,\leftrightarrow,\to,\leq)$ is the extension of Boolean lattice $BL=(X,1,0,\neg,\wedge,\vee,\leq)$ by defining the new operations ($\leftrightarrow$ and $\to$) with the axioms below:
@@ 5     $=x[x\mapsto y]\leftrightarrow x$ Selected replacement
6     $=y\leftrightarrow x$ Result of the replacement
7 $x\leftrightarrow y\leq y\leftrightarrow x$ 2–6: Transitivity of $=$ and $\leq$
@@ 2     $\leq(x\leftrightarrow y)[y\mapsto z]$ Replacement property of $\leftrightarrow$ (Axiom 1)
3     $=(x\leftrightarrow z)$ Result of the replacement
@@
Theorem 4.3.4 (Leibniz’s replacement in IRLA).
@@ 9 $x=y$ 5,8: Antisymmetry of $\leq$
Thus we have $x\leftrightarrow y=1\Rightarrow x=y$, and combining this with $x=y\Rightarrow x\leftrightarrow y=1$ (Axiom 2) we get $x\leftrightarrow y=1\Leftrightarrow x=y$
@@
By definition of $\to$ (Axiom 3), $x\to y=1\Leftrightarrow x\wedge y\leftrightarrow x=1$, and $\leq$ is defined by $x\leq y\Leftrightarrow x\wedge y=x$. In terms of the above theorem, $x\wedge y\leftrightarrow x=1\Leftrightarrow x\wedge y=x$, so $x\to y=1\Leftrightarrow x\leq y$. ∎
In terms of Theorem 4.3.5 and Corollary 4.3.1, an IRLA theorem of the form $s\leftrightarrow t=1$ or $s\to t=1$ follows immediately from a corresponding theorem of Boolean lattice (BL) of the form $s=t$ or $s\leq t$, so the latter is also a theorem in IRLA. Some examples of such theorems of IRLA and BL are listed in Table 3.
Table 3: Some theorems of IRLA and BL
Name Form only in IRLA Form in BL and IRLA
@@
Theorem 4.3.6 (Bi-implication introduction / Antisymmetry of implication in IRLA).
@@
1     $=(x\leftrightarrow x\wedge y)\wedge(x\wedge y\leftrightarrow y)$ Definition of $\to$
@@ 2     $\leq x\leftrightarrow y$ Transitivity of $\leftrightarrow$
3 $(x\to y)\wedge(y\to x)\leq x\leftrightarrow y$ 1–2: Transitivity of $=$ and $\leq$
Theorem 4.3.7 (Bi-implication elimination in IRLA).
$x\leftrightarrow y\leq x\to y,\>x\leftrightarrow y\leq y\to x.$
Proof.
$x\leftrightarrow y$
1     $=(x\leftrightarrow y)\wedge 1$ Identity element
2     $=(x\leftrightarrow y)\wedge(x\to x)$ Reflexivity of $\to$ ($x\leftrightarrow x=1$)
3     $\leq(x\to x)[x\mapsto y]$ Replacement property of $\leftrightarrow$ (Axiom 1)
4     $=x\to x[x\mapsto y]$ Selected replacement
5     $=x\to y$ Result of the replacement
6 $x\leftrightarrow y\leq x\to y$ 1–5: Transitivity of $=$ and $\leq$
The second is obtained by symmetry of $\leftrightarrow$. ∎
Theorem 4.3.8 (Bi-implication by implication in IRLA).
$x\leftrightarrow y=(x\to y)\wedge(y\to x).$
Proof.
1 $x\leftrightarrow y\leq x\to y$ Bi-implication elimination
2 $x\leftrightarrow y\leq y\to x$ Bi-implication elimination
3 $x\leftrightarrow y\leq(x\to y)\wedge(y\to x)$ 1,2: One-to-two of $\leq$
4 $(x\to y)\wedge(y\to x)\leq x\leftrightarrow y$ Bi-implication introduction
5 $x\leftrightarrow y=(x\to y)\wedge(y\to x)$ 3,4: Antisymmetry of $\leq$
@@ 1     $=(x\wedge y\leftrightarrow x)\wedge(y\wedge z\leftrightarrow y)$ Definition of $\to$
3     $\leq(x\wedge y\leftrightarrow x)\wedge(x\wedge(y\wedge z)\leftrightarrow x)$ Replacement $[y\mapsto y\wedge z]$
@@ 4     $=(x\wedge y\leftrightarrow x)\wedge((x\wedge y)\wedge z\leftrightarrow x)$ Associativity
5     $\leq x\wedge z\leftrightarrow x$ Replacement $[x\wedge y\mapsto x]$
@@ 1     $=(x\wedge z\leftrightarrow z)\wedge(y\wedge z\to z)$ Definition of $\to$
2     $\leq x\wedge y\wedge z\leftrightarrow z$ Replacement $[z\mapsto y\wedge z]$
@@ 4 $(z\to x)\wedge(z\to y)\leq z\to x\wedge y$ 1–3: Transitivity of $=$ and $\leq$
5     $=(z\to x\wedge y)\wedge(x\wedge y\to x)$ Conjunction elimination ($x\wedge y\to x=1$)
@@ 10 $(z\to x)\wedge(z\to y)=z\to x\wedge y$ 4,9: Antisymmetry of $\leq$
Theorem 4.3.11 (Modus ponens in IRLA).
$(x\to y)\wedge x\leq y.$
Proof.
$(x\to y)\wedge x$
1     $=(x\leftrightarrow x\wedge y)\wedge x$ Definition of $\to$
2     $\leq x\wedge y$ Replacement property of $\leftrightarrow$ ($[x\mapsto x\wedge y]$)
3     $\leq y$ Conjunction elimination
4 $(x\to y)\wedge x\leq y$ 1–3: Transitivity of $=$ and $\leq$
Corollary 4.3.2 (Upper bound of implication).
@@ 4 $(x\to y)\wedge x\leq y$ 1–3: Transitivity of $=$ and $\leq$
In BL, $x\wedge y\leq z\Leftrightarrow x\leq\neg y\vee z$, so modus ponens $(x\to y)\wedge x\leq y$ is equivalent to the upper bound of implication $x\to y\leq\neg x\vee y$ in IRLA, hence the latter follows from the above theorem. ∎
@@
It follows from boundedness $0\leq 1\to 0$ and $1\to 0\leq\neg 1\vee 0=0\vee 0=0$ by Corollary 4.3.2. ∎
@@ 4.4 Soundness and completeness
By definition (cf. Subsection 4.1), the T-set of IRL formula schema $F(\phi,\psi,...)$ is the IRLA term $t(x,y,...)$; and $\vDash\phi$ is the same as $x=1$, $\vDash\phi\leftrightarrow\psi$ is the same as $x=y$ or $x\leftrightarrow y=1$, and $\vDash\phi\to\psi$ is the same as $x\leq y$ or $x\to y=1$.
@@
For each axiom schema $F(\phi,\psi,...)$ of IRL, verify $t(x,y,...)=1$ in IRLA. For example, $\phi\leftrightarrow\phi$ is a tautology since $x\leftrightarrow x=1$ in IRLA. ∎
@@
All the three primitive rules (except for the meta rule of conditional proof) of IRL, i.e. Conjunction introduction $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$, Conjunction elimination $\phi\wedge\psi\Rightarrow\phi$ and Replacement property $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$, are tautologousness-preserving, that is, if their premises are tautologies, so are their conclusions.